OpenStudy (anonymous):

On the graph of the equation 5x − 2y = 20, what is the value of the y-intercept? −10 10 −4 4

OpenStudy (gorv):

to calculate y intercept put x =0

OpenStudy (anonymous):

@sidona005 can you help?

OpenStudy (anonymous):

When y-intercepts, x =0 5(0) - 2y = 20 sooooo 5(0) - 2y = 20 0-2y = 20 -2y=20 y=-10

OpenStudy (anonymous):

thanks so much your alot of help